OpenStudy (anonymous):

I have to find the number of ham subs with any combination of toppings or no toppings at all. The toppings are lettuce, onions, peppers, and olives Can someone please help???

OpenStudy (ybarrap):

You can either have lettuce or no lettuce. That's 2 possibilities for the 1st topping You can either have onions or no onions. That's 2 possibilities for the 2nd topping and so on The number of combinations of lettuce, onions, peppers and olives then is $$ \large 2\times 2\times2\times 2=2^4 $$ Does this make sense?